find the values for x, y, and z such the matrix below is skew symmetric.

0 x 3

2 y -1

z 1 0


1
Expert's answer
2022-05-17T09:29:07-0400
"A=\begin{pmatrix}
 0 & x & 3\\
 2 & y & -1 \\
 z & 1 & 0 \\
\end{pmatrix}"

Given matrix is skew symmetric matrix:


"A'=-A"

"\begin{pmatrix}
 0 & 2 & z\\
 x & y & 1 \\
 3 & -1 & 0 \\
\end{pmatrix}=\begin{pmatrix}
 0 & -x & -3\\
 -2 & -y & 1 \\
 -z & -1 & 0 \\
\end{pmatrix}"

On comparing we get:


"x=-2, y=0, z=-3"